OpenStudy (anonymous):

Simplify the complex fraction [a/b - b/a][a - a^2/(a+b)]

OpenStudy (anonymous):

a - b

OpenStudy (anonymous):

over b?

OpenStudy (anonymous):

I got (a-b)/b

OpenStudy (anonymous):

[(a^2 - b^2)/ab] [(a^2 + ab - a^2)/(a+b) = [(a^2 - b^2)/ab][ab/a + b] =(a^2 - b^2) / (a+b)=a -b

OpenStudy (anonymous):

alright thanks guys!

OpenStudy (anonymous):

All the best outcastlilgurl

OpenStudy (anonymous):

isnt a-a^2=a(1-a), not a^2+ab-a^2

OpenStudy (anonymous):

it is a^2 over a+b, not a - a^2 over a+b

OpenStudy (anonymous):

Ohhhhh i see, i did a-a^2 all over b. gotcha

OpenStudy (radar):

Is the answer a-b?

OpenStudy (radar):

(a/b - b/a)=\[(a ^{2}-b ^{2})/ab\] that would be the first product

OpenStudy (radar):

\[a-a ^{2}/(a+b)=(a ^{2}+ab-a ^{2})/(a+b)\] is the second product.

OpenStudy (radar):

Now multiply them: First factor of the product: Second factor of the product\[(a ^{2}-b ^{2})/ab \times(a ^{2}+ab-a ^{2})/(a+b)\] Simplifying\[[(a+b)(a-b)]/ab \times ab/(a+b)\] Further simplifying:\[a-b\] remaining terms cancel out